OpenStudy (anonymous):

Detremine the root mean square speed of CO2 molecules that have an average kenetic energy of 4.21 X 10^-21

OpenStudy (aroub):

kinetic energy=1/2mv^2

OpenStudy (aroub):

m is the molar mass, and v is the speed.

OpenStudy (aroub):

First step: Find the molar mass of CO2

OpenStudy (aroub):

Then you just substitute & simplify :)

OpenStudy (anonymous):

this is what I got when I worked it out but its saying its wrong. 9.26399975 X 10^-20

OpenStudy (aroub):

What did you do exactly?

OpenStudy (aaronq):

\(\Large v_{rms}=\sqrt{\dfrac{2KE}{m_{CO_2}}}\) m= mass of onle molecule